Anna and Doug are in the market for a new house. The maximum payment they can afford is $700 per month. Of this payment, property taxes and homeowner’s insurance amount to $150 per month. If the interest rate on the mortgage is 4.5% per year, how much house can Anna and Doug afford to finance? The duration of the mortgage loan is 30 years (360 months).
Answer #1

Duration of the loan = 360 months

Payment per month = 700-150 = 550

Interest rate = 4.5%

Monthly interest rate = 4.5/12

Now we consider A=550 its a series of annuity payments up till 360 months

Now to calculate the present value of this annuity we use (P/A,4.5%/12,360)

=550(P/A,4.5%/12,360)

=550x197.361159

$108548.637